OpenStudy (anonymous):

What is the solution set of |x – 2| + 4 = 3?

Parth (parthkohli):

An absolute value equal to a negative number.

OpenStudy (zugzwang):

What he said^

Parth (parthkohli):

\[|x - 2| = -1\]Hmm

OpenStudy (zugzwang):

No... See, that absolute value part? It's never going to be negative. It could be zero, but never negative. So you add something non-negative to 4 And you get 3 I don't want to live on this planet anymore XD

OpenStudy (ryan123345):

its not 3 or negative 3....

OpenStudy (anonymous):

Zugzwang, I can help you with that...;)

OpenStudy (zugzwang):

Well, I shall wait until you make good on that :P

OpenStudy (zugzwang):

It's an important and popular fact that for any real number a, |a| ≥ 0

OpenStudy (ryan123345):

i dont think there is any solutions...

OpenStudy (zugzwang):

There aren't... given that 'important and popular fact', |x - 2| + 4 = 3 |x - 2| + 4 - 4 = 3 - 4 |x - 2| = -1 And like the fact says, this cannot be, as the absolute value of any real number is never going to be negative.

OpenStudy (ryan123345):

So it would be ''no solutions''

Parth (parthkohli):

Yes.

OpenStudy (zugzwang):

What he said^

Parth (parthkohli):

What he said^

OpenStudy (ryan123345):

lol.

OpenStudy (anonymous):

Alt+f4

OpenStudy (zugzwang):

You are about to close 3 open tabs... bla bla bla

OpenStudy (anonymous):

=)

OpenStudy (ryan123345):

you should click start that shut down to see what happends c:

OpenStudy (ryan123345):

Than*

OpenStudy (ryan123345):

you should click start than shut down to see what happends c:

OpenStudy (zugzwang):

then*

OpenStudy (ryan123345):

Thank you very much c:

OpenStudy (ryan123345):

lol